OpenStudy (anonymous):

find a linear function with slope -3 that contain (-1,4) graph the function

OpenStudy (anonymous):

You've got \[(x,y)=(-1,4)\] And the linear function: \[f(x)=a\cdot x+b\] Find b: \[f(-1)=-3\cdot -1+b\]

OpenStudy (anonymous):

\[f(-1)=4\] of course

OpenStudy (anonymous):

When you've found b, you just insert a and b into the linear function and let x be your variable.

OpenStudy (anonymous):

you could use the equation m=(y-y1)/(x-x1) m would be the slope of your function y1 the y component of the coordinate and x1 the x component of the coordinate therefore, -3=(y-4)/(x-(-1)) -3=(y-4)/(x+1) cross multiplying the equation gives us y-4=-3(x+1) y-4=-3x-3 y=-3x-3+4 y=-3x+1 f(x)=-3x+1,voila, you got your linear funciton since this is is a linear function, you could draw the graph using only the x and y intercepts which can be found by i.)letting y=0 for the x intercept thus, 0=-3x+1 3x=1 x=1/3 therefore your graphed function would cut the x axis at the point (1/3,0) ii.)letting x=o for the y intercept thus, y=-3(0)+1 y=1 therefore your graphed function should cut the y axis at the point (0,1) HINT: Just plot the points (1/3,0) and (0,1) on the x-y planes and join them using a straight line and you will have graphed your function

OpenStudy (anonymous):

if there's anything you haven't quite understood just let me know